OpenStudy (anonymous):

MEDAL!~ Verify my answer please. What is the area of an isosceles right triangle if the side opposite the right angle is 4 in? my answer: approx 5.66in^2

OpenStudy (triciaal):

side opposite the right angle is the hypotenuse = 4 isosceles triangle 2 sides equal let this = a Pythagorean theorem a^2 + a^2 = 4^2 2a^2 = 16 so a =rt8 Area = 1/2 base * height = 1/2 *rt8 *rt8 rt8 * rt8 = 8 I got area = 4

OpenStudy (anonymous):

@triciaal I probably made a mistake. Uhm, can't I use this solution? (see attachment)

OpenStudy (anonymous):

I'm confused as to when I should use the pythagorean theorem and the trigo functions >.< @triciaal

OpenStudy (triciaal):

in your diagram a and c are switched c is the hypotenuse opposite the rt angle

OpenStudy (anonymous):

Ah, that makes sense but using trigo functions is a no-no? D:> there's a slight difference between your answer and mine @triciaal